OpenStudy (anonymous):

HELP. x over 2x-4-2 = 1 over x-2

OpenStudy (anonymous):

is the left side correct - it has two numbers -4 and -2

OpenStudy (anonymous):

ya its x over 2x-4 then -2 = 1 over x-2

OpenStudy (anonymous):

xo you know how to do it?

OpenStudy (anonymous):

oh ok - sorry so x / (2x -4) - 2 = 1 / (x-2) multiply thru by 2x-4 : x - 2(2x-4) = 2 x - 4x + 8 = 2 3x = 6 x = 2

OpenStudy (anonymous):

sorry i was so long but i have to type left handed as i broke my right arm

OpenStudy (anonymous):

no worries sorry bout that thanks so much

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

wait its suppost to come out to no solution i dont know whats wrong